# Investigation of the effects of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere using spatial dynamic mode decomposition

### v1

@@ 3 Results
Figure 6: Eigenvalues of (a) $\widetilde{\boldsymbol{A}}$ and (b) $\boldsymbol{A}$ from spatial DMD. The black dashed line in (a) is the unit circle.
Table 2: Decay rates $r^{x}_{k}$, per diameter, and wavelengths $\lambda^{x}_{k}$, in diameters, of the first three DMD modes in the wakes of the REF, PRS and SHS spheres.
DMD Mode, $k$ 1 2 3
REF $r^{x}_{k}$0.07 -0.08 -0.10
$\lambda^{x}_{k}$ 4.48 2.27 1.95
PRS $r^{x}_{k}$0.06 -0.07 -0.08
$\lambda^{x}_{k}$ 4.52 2.22 2.40
SHS $r^{x}_{k}$0.06 -0.11 0.13
$\lambda^{x}_{k}$ 3.08 9.70 22.67
All DMD modes resulting from the spatial dynamic mode decomposition have negative real parts, as shown in figure 6b, meaning that they all decay in the streamwise direction. This is expected, as turbulent flows are typically stable and do not feature convective instabilities. The first two DMD modes of the REF and PRS spheres have very similar decay rates and wavenumbers, while there is a significant difference in the decay rates and wavenumbers of their third DMD modes. This suggests that the two DMD modes with the slowest decay rates are unaffected by the addition of the pores, while the pores have some effect on the dynamics that decay more quickly in the streamwise direction. The decay rates and wavenumbers of the first three DMD modes in the wake of the SHS sphere are significantly different from those of the untreated spheres, suggesting that the addition of superhydrophobic surface treatment has a significant effect on the dynamics in the near wake of the sphere. This is consistent with the leading DMD modes of the proper orthogonal decomposition, compared in [38], which are very similar for the untreated spheres but radically different for the SHS sphere.
Figure 7: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the first DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The first DMD modes are qualitatively similar between the REF and PRS spheres, as shown in figures 7a and 7b, respectively. The streamwise components of the real part of the first DMD modes for each sphere exhibit antisymmetry across $y=0$, with regions of alternating sign in the streamwise direction, as shown in figures 7a(i) and 7b(i), respectively. The corresponding transverse components are symmetric across $y=0$, also with regions of alternating sign in the streamwise direction, as shown in figures 7a(ii) and 8b(ii), respectively. This behaviour is reflected in the imaginary parts of the streamwise components, as shown in figure 7a(iii) and 7b(iii), respectively, and the transverse components, shown in figure 7a(iv) and 7b(iv), respectively, which show the same structures propagating downstream. In addition to the qualitative similarity, the slowest decaying DMD modes of the REF and PRS spheres have comparable wavelengths of $4.48D$ and $4.52D$, respectively. The decay rate of the first DMD mode of the PRS sphere decays slightly more slowly than that of the REF sphere, at $-0.06D^{-1}$ and $-0.07D^{-1}$, respectively. Although qualitatively similar in structure to the first DMD modes of the REF and PRS spheres and decaying at a similar rate of $-0.06D^{-1}$, the first DMD mode of the SHS sphere, shown in figure 7c, has a significantly shorter wavelength of $3.08D$, which is reflected in the reduced length of the structures in the SHS mode compared to the REF and PRS modes.
Figure 8: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the second DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The second DMD modes of the REF and PRS spheres, shown in figures 8a and 8b, respectively, are qualitatively similar to each other and have comparable decay rates, $-0.08D^{-1}$ and $-0.07D^{-1}$, and wavelengths, $2.27D^{-1}$ and $2.22D^{-1}$, respectively. Additionally, the structure of these modes is similar to that of the first mode but with the length of the structures in the streamwise direction reduced, which is reflected in the wavelengths being decreased by a factor of two compared to those of the first modes of each sphere. As with the first DMD modes of these spheres, the imaginary part of the second DMD modes, the streamwise and transverse components of which are shown in figures 8a(iii) and 8b(iii), and figures 8a(iv) and 8b(iv), respectively, represent the propagation of the real parts, the streamwise and transverse components of which are shown in figures 8a(i) and 8b(i), and figures 8a(ii) and 8b(ii), respectively, in the streamwise direction. The second DMD mode in the SHS case, shown in figure 8c, is substantially different from those of the untreated spheres. This is particularly evident when comparing the streamwise component of the second DMD mode, where the positive region directly behind the sphere, shown in figure 8c(i), becomes substantially larger and moves off centre as the negative region at $y/D=0.5$ is replaced with asymmetric negative regions at $y/D=\pm 0.5$, shown in figure 8c(iii).
Figure 9: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the third DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The third DMD modes of the REF and PRS spheres are also qualitatively similar, although the third PRS mode, shown in figure 9b, is significantly less symmetric than the third REF mode, shown in figure 9a. The asymmetry is likely due to the spherical distribution of pores on the PRS sphere, as the surface of the sphere on either side of the plane measured may not have exactly the same distribution of pores, resulting in an uneven boundary layer across the sphere. In addition to the introduction of asymmetry, the PRS mode has a lower decay rate of $-0.08D^{-1}$ compared to $-0.10D^{-1}$ for the REF sphere, and a longer wavelength of $2.40D$ compared to $1.95D$. This is reflected in the size of the structures, which are larger in the PRS case than in the REF case. As in the second DMD mode, the third DMD mode of the SHS sphere, shown in figure 9c, exhibits large structures that have substantial transverse variation in place of the streamwise propagation seen in the modes of the untreated spheres.
Figure 10: (i) Real and (ii) imaginary out-of-plane vorticity of the first DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The periodicity in the streamwise direction of the first DMD modes in the wake of each sphere is seen clearly in their out-of-plane vorticity, shown in figure 10. In addition to the reduced wavelength of the structures in the wake of the SHS sphere seen in the velocity components, the vorticity contour, shown in figure 10c, shows a planar asymmetry where the untreated spheres are planar symmetric, as shown in figures 10a and 10b. The growth of the wake can also be seen in these structures, whose transverse extent increases in the streamwise direction.
Figure 11: (i) Real and (ii) imaginary out-of-plane vorticity of the second DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The out-of-plane vorticity contours of the second DMD mode in the wakes of the REF and PRS spheres, shown in figures 11a and 11b, respectively, show similar structures to the first DMD modes but with reduced wavelength. This is consistent with the velocity components of these modes, which were also qualitatively similar. The vortical structures in the second DMD mode of the SHS sphere, shown in figure 11c, are substantially different from the untreated spheres. However, the positive vorticity directly behind the sphere is bounded by thinner negative regions with the sign of these regions changing after $x/D=3$, which is similar to the structures seen in the first and second DMD modes of the untreated spheres. Unlike the REF and PRS cases, the imaginary part of the second SHS mode, shown in figure 11c(ii), does not resemble a streamwise translation of its real part, shown in figure 11c(i). Although the large positive region in the imaginary part is bounded by negative regions, the pattern is heavily biased in the transverse direction, and the streamwise extent of the positive region is significantly longer than in the real part. An additional negative region is also present directly behind the sphere, within the mean recirculation region.
Figure 12: (i) Real and (ii) imaginary out-of-plane vorticity of the third DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The out-of-plane vorticity of the third DMD mode in the REF sphere, shown in figure 12a, is planar antisymmetric with structures directly behind the sphere bounded by structures of opposing sign. As with the velocity components, shown in figure 9a, the intensity of the out-of-plane vorticity decays in the streamwise direction. The out-of-plane vorticity of the third DMD mode in the PRS sphere, shown in figure 12b, shows the same structures elongated in the streamwise direction, with the planar asymmetry seen in the velocity, shown in figure 9b, components reflected in the out-of-plane vorticity. The out-of-plane vorticity of the third DMD mode of the SHS case, shown in figure 12c, shows no streamwise periodicity, as was the case with the velocity components. The imaginary part of the out-of-plane vorticity, shown in figure 12c(ii), appears to represent a rotation of the structures in the real part, shown in figure 12c(i), around the streamwise axis, instead of a propagation in the streamwise direction, as seen in the third DMD modes of the untreated spheres.
@@ 4 Discussion
The shapes of the first and second DMD modes of the REF and PRS spheres, and the first mode of the SHS sphere, are similar and consistent with the flapping of the wake. The second DMD modes of the REF and PRS cases are harmonics of the corresponding first DMD modes, as the wavelengths of the second DMD modes are approximately half those of the first DMD modes. The third DMD modes of the REF and PRS cases represent a pulsating motion, as alternating positive and negative streamwise velocity fluctuations propagate downstream. The first dynamic mode of the SHS sphere represents a flapping motion of the wake, as is the case for the untreated spheres. However, the wavelength and decay rate of this fluid motion are significantly altered by the superhydrophobic surface treatment.
The second and third DMD modes in the SHS case are substantially different from those of the REF and PRS cases. There is a more significant change in decay rate between the first DMD mode and the second and third DMD modes in the SHS case than in the case of the untreated spheres. Additionally, the wavelengths of the second and third DMD modes in the SHS case are significantly longer than the measurement domain, which is $4.7D$ in the streamwise direction. This suggests that these DMD modes may not be periodic in the streamwise direction, but they extend far enough downstream of the sphere that this cannot be verified with the current measurement domain. This is reflected in the DMD mode shapes, which do not exhibit the streamwise patterns observed in the other DMD modes. Despite the substantial distortion, the basic vortical structure of the real part of the second DMD mode of the SHS case is similar to that of the first SHS mode and the first and second DMD modes in the REF and PRS cases. The third DMD mode of the SHS sphere shows significantly more substantial three-dimensional effects than the third DMD modes of the REF and PRS spheres, with its real and imaginary parts representing a rotation around the streamwise axis rather than a propagation in the streamwise direction, as seen in the DMD modes of the untreated spheres.
While the leading DMD modes in the wake of the untreated spheres all exhibit a degree of planar symmetry, only the first DMD mode in the SHS case shows any degree of planar symmetry. The second and third DMD modes in the wake of the SHS sphere show more significant three-dimensionality than those of the untreated spheres. The reduction in planar symmetry in the DMD modes is most likely due to the changing shape of the plastron as it is subjected to pressure fluctuations across the sphere, as observed in the wakes of spheres with superhydrophobic surface treatment settling under gravity [14]. As the plastron is deformed by the fluid flow, the changing shape of the plastron affects the flow in turn, resulting in a feedback loop like that proposed in [47], which makes the flow significantly more complex dynamically than that in the wake of the untreated spheres. This added complexity, which would vary significantly with different geometries and flow conditions, is consistent with the highly variable drag effects seen in previous investigations of superhydrophobic surface treatments [19, 21].

### v2

@@ 3 Results
Table 3: Magnitudes $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ and angles $\angle\mathcal{R}_{+,-}^{\mathbb{C}}$ of the cross-domain complex correlation coefficients of the DMD modes of each sphere.
REF PRS SHS
$k$ $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ $\angle\mathcal{R}_{+,-}^{\mathbb{C}}$ $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ $\angle\mathcal{R}_{+,-}^{\mathbb{C}}$ $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ $\angle\mathcal{R}_{+,-}^{\mathbb{C}}$
[…]
The modes of the REF sphere have consistently high $|\mathcal{R}_{+,-}^{\mathbb{C}}|$, with the lowest being 0.95 for the fifth mode, suggesting that these modes are highly symmetric/antisymmetric across the centreline. The PRS modes also have high $|\mathcal{R}_{+,-}^{\mathbb{C}}|$, although the values for the first and fifth modes are noticeably lower than the rest at 0.89 and 0.80, respectively. The values of $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ for the SHS sphere are more varied, ranging from 0.77 for the second mode to 0.96 for the fourth mode. This suggests that the symmetry/antisymmetry of these modes is affected by the addition of superhydrophobic surface treatment.
Figure 6: DMD eigenvalues of (a) $\widetilde{\boldsymbol{A}}$, (b) $\boldsymbol{A}$, and (c) energy shares $s_{k}$ for each sphere. Hollow symbols denote conjugate pairs. The dashed line in (a) represents the unit circle.
The $\widetilde{\mu}_{k}$ for all three spheres sit inside the unit circle, as shown in figure 6a, and thus the corresponding $\mu_{k}$ have negative real parts, as shown in figure 6b. This is consistent with the turbulence decaying downstream of the sphere. The energy shares of the modes are similar between the REF and PRS spheres, suggesting that the effect of the pores on the dynamic modes is relatively minor compared to the effect of the superhydrophobic surface treatment, as shown in figure 6c.
Table 4: Energy shares $s_{k}$, decay rates $r_{k}^{x}$ and wavelengths $\lambda_{k}^{x}$ of the DMD modes of each sphere.
REF PRS SHS
$k$ $s_{k}$ [%] $r_{k}^{x}D$ $\lambda_{k}^{x}/D$ $s_{k}$ [%] $r_{k}^{x}D$ $\lambda_{k}^{x}/D$ $s_{k}$ [%] $r_{k}^{x}D$ $\lambda_{k}^{x}/D$
1 22.14 -0.02 5.38 21.03 -0.02 5.54 15.96 -0.05 2.99
2 19.88 -0.04 2.64 18.76 -0.03 2.70 14.82 -0.03 4.38
3 13.91 -0.11 1.58 15.65 -0.10 1.72 13.56 -0.07 1.95
4 10.66 -0.19 0.99 11.78 -0.14 1.06 10.03 -0.11 1.18
5 10.24 -0.09 3.86 11.12 -0.07 3.84 23.67 -0.10 65.38
6 23.17 -0.11 53.89 21.65 -0.09 36.33 21.96 -0.10 10.44
Both the REF and PRS spheres feature five DMD modes with wavelengths shorter than twice the measurement domain and one DMD mode with a wavelength significantly longer than the measurement domain. Each of the first five modes of these spheres exhibit similar energy shares, decay rates and wavelengths, as shown in table 4. Only four of the SHS modes have wavelengths shorter than twice the measurement domain, with the fifth mode having a wavelength significantly above this threshold and the sixth mode being only slightly above, at 65.38DD and 10.44DD, respectively. The sixth REF and PRS modes each have the largest energy share, despite the wavelength of these modes being significantly longer than the measurement domain, at 53.89DD and 36.33DD, respectively. This is also the case for the fifth and sixth SHS modes. Despite all spheres exhibiting modes with long wavelengths, the values of these wavelengths vary significantly between spheres.
Figure 7: Correlation maps betwwen the modes of the (a) REF and PRS spheres, (b) REF and SHS spheres, and (c) PRS and SHS spheres.
The similarity in the eigenvalues of the first five REF and PRS modes is reflected in the degree of diagonality of the cross-correlation matrix, as shown in figure 7a. The cross-correlation between the SHS and untreated spheres is less strongly diagonal, as shown in figures 7b and 7c, for the REF and PRS spheres, respectively. However, there are some consistent features between these cross-correlation maps which suggest that some SHS modes are related to the modes of the untreated spheres. The first and second SHS modes best match the second and first REF and PRS modes, respectively. The third and fourth modes of the SHS are most highly correlated with the corresponding REF and PRS modes. This is also consistent with the first modes of the REF and PRS spheres having a longer wavelength than the corresponding second modes, while the second SHS mode has a longer wavelength than the first SHS mode. While the energy share of the first two SHS modes is substantially lower than the leading modes of the untreated spheres, at approximately 15% each compared with 20% each, the first and second modes have comparable energy shares for each sphere. The third and fourth modes have similar wavelengths and energy shares across all three spheres.
Figure 8: Normalised (i) streamwise and (ii) transverse velocity components, and corresponding (iii) out-of-plane vorticity for the six (a)-(f) DMD modes of the REF sphere.
Figure 9: Normalised (i) streamwise and (ii) transverse velocity components, and corresponding (iii) out-of-plane vorticity for the six (a)-(f) DMD modes of the PRS sphere.
Figure 10: Normalised (i) streamwise and (ii) transverse velocity components, and corresponding (iii) out-of-plane vorticity for the six (a)-(f) DMD modes of the SHS sphere.
The first REF mode has an energy share of 22.1422.14%, a decay rate of 0.02/D-0.02/D, and a wavelength of $5.38D$. The first PRS mode has similar values, with an energy share of 21.0321.03%, a decay rate of 0.02/D-0.02/D, and a wavelength of $5.54D$. The first REF mode is more strongly symmetric/antisymmetric than the PRS mode, as shown by their $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ values of 0.96 and 0.89, respectively. The REF mode, shown in figure 8a, features regions of streamwise and transverse fluctuations of alternating sign reminiscent of the first two POD modes in [40] and first Hilbert POD mode in [51], which represent a flapping in the wake. The PRS mode, shown in figure 9a, features similar structures, although these are more distorted relative to the first POD mode [40]. The highest intensity in the streamwise component of the REF mode, shown in figure 8a(i), remain centred at $r/D\approx 0.5$ for the entire measurement domain. In contrast, the structures close to the sphere in the streamwise component of the PRS mode, shown in figure 9a(i), have the highest intensity closer to the centreline. Further downstream, the structures in the streamwise component of the PRS mode shift away from the centreline to match those of the REF mode. The shift in the structures in the PRS mode relative to the REF mode is reflected in the out-of-plane vorticity, which features longer horizontal structures on the REF mode, as shown in figure 8a(iii), and shorter slanted structures in the PRS case, shown in figure 9a(iii). From the eigenvalues, the closest match to these modes is the second mode of the SHS sphere, which has an energy share of 14.82%, a decay rate of 0.03/D-0.03/D, and a wavelength of $4.38D$. In addition to having a shorter wavelength than the untreated spheres, the structure of this mode is distinct from that of the corresponding modes of the untreated spheres. Directly behind the sphere, the structures in both the streamwise and transverse components, shown in figures 10b(i) and 10b(ii), respectively, curve towards the centreline before growing larger and shifting in the transverse direction further downstream. While both of the untreated spheres feature structures which are completely separated in the streamwise direction, the structures of the SHS mode overlap, with the downstream structure extending over the upstream structure around $x/D=2$. Further downstream, at $x/D=4$, the structures no longer overlap, and the shape of the out-of-plane vorticity changes substantially from the approximately horizontal structures upstream, as shown in figure 10b(iii).
The second REF mode, shown in figure 8b, has an energy share of 19.8819.88%, a decay rate of 0.04/D-0.04/D, and a wavelength of $2.64D$. The second PRS mode, shown in figure 9b, has an energy share of 18.7618.76%, a decay rate of 0.03/D-0.03/D, and a wavelength of $2.70D$. The wavelengths of both modes are approximately half of the corresponding first modes, and their structures are similarly comparable. This is consistent with the sixth and eighth POD modes and third Hilbert POD mode of REF sphere in [51]. As in the case of the first mode, the REF mode features horizontal structures, while the structures in the PRS sphere are relatively deformed. This is most evident in the out-of-plane vorticity, which is uniform in the streamwise direction in the REF mode, as shown in figure 8b(iii), while the structures of the out-of-plane vorticity of the PRS mode, shown in figure 9b(iii), change in size and orientation as they advect downstream. The first SHS mode is the closest match to the second modes of the untreated spheres, with an energy share of 15.96%, a decay rate of 0.05/D-0.05/D, and a wavelength of $2.99D$. Unlike the second SHS mode, which was closest to the first modes of the untreated spheres, this mode has a longer wavelength than the second modes of the untreated spheres. While the downstream structures of the second SHS mode are similar to the first SHS mode, particularly in the out-of-plane vorticity, shown in figure 10a(iii), the structure directly behind the sphere is significantly altered. In addition to the large structures which alternate in sign downstream, smaller structures are present for $x/D<2$. In the streamwise component, shown in figure 10a(i), there are two small regions of high intensity at $r/D=0.5$ in addition to the larger structures present in the corresponding modes of the untreated spheres. The corresponding region in the transverse component, shown in figure 10a(ii), features a low-intensity positive region that is not present in the modes of the untreated spheres. Additionally, the negative region closest to the sphere features a vertical protrusion on the upstream side, which is more intense than the corresponding structure in the second PRS mode. These structures are reflected in the out-of-plane vorticity of the first SHS mode, shown in figure 10a(iii), which features small staggered high-intensity regions close to the sphere and large diagonal structures further downstream, the latter of which are more reminiscent of the structures in the second SHS mode.
The third modes of each sphere are closest matches to one another, with energy shares of $12.91\%$, $15.65\%$, and $13.56\%$, decay rates of 0.11/D-0.11/D, 0.10/D-0.10/D, and 0.07/D-0.07/D, and wavelengths of $1.58D$, $1.72D$, and $1.95D$, for the REF, PRS, and SHS spheres, respectively. All three modes, shown in figures 8c, 9c and 10c, respectively, feature regions of alternating sign in their streamwise component, which are located around $r/D=0.5$ close to the sphere, and move closer to the centreline as they advect downstream. The transverse components exhibit similar behaviour, although the these structures have an aspect ratio close to unity, while the structures in the streamwise velocity are longer in the streamwise direction, particularly close to the sphere. The out-of-plane vorticity of these modes consists of uniform structures which fade in intensity as they travel downstream, which are most intense at $r/D=0.5$ and $x/D<1$. Unlike the first and second modes, there is no obvious change in the mode structures from either the addition of pores or superhydrophobic surface treatment. Additionally, this mode does not clearly relate to the POD modes in [40], which were similar for the untreated spheres, but significantly altered in the SHS case. The fourth modes of each sphere are mutually best matches, with energy shares of $0.24\%, 11$$1.12\%$, and $10.03\%$, decay rates of 0.19/D-0.19/D, 0.14/D-0.14/D, and 0.11/D-0.11/D, and wavelengths of $0.99D$, $1.06D$, and $1.18D$, for the REF, PRS, and SHS spheres, respectively. These modes, shown in figures 8d, 9d and 10d, are similar in shape to the third modes of each sphere, but contain smaller structures corresponding to the shorter wavelengths of these modes.
The fifth modes of the REF and PRS spheres have similar energy shares of $10.24\%$ and $11.12\%$, respectively. Their eigenvalues are also similar, with decay rates of 0.09/D-0.09/D and 0.07/D-0.07/D, and wavelengths of $3.86D$ and $3.84D$, respectively. The modes are similar in structure, although the PRS mode, shown in figure 9e, is phase-shifted relative to the REF mode, shown in figure 8e. The intensity of these modes drops off sharply after $x/D\approx 3$. This corresponds to the structural change from parallel regions of opposing sign stacked in the radial direction to regions of alternating sign in the streamwise direction.
The sixth modes of both the REF and PRS spheres are characterised by wavelengths of $53.89D$ and $36.33D$, respectively, which are significantly longer than the measurement domain. Despite this, these modes have the highest energy share for their respective spheres, at $23.17\%$ and $21.65\%$. There is only a single structure in the velocity components of these spheres, as shown in figure 8f and 9f, respectively. However, the out-of-plane vorticity of these modes, and particularly the PRS mode, resembles the structures in the first modes of the untreated spheres, but extends beyond the measurement domain. The fifth and sixth SHS modes, shown in figure 10e and 10f, have wavelengths of $65.38D$ and $10.44D$, respectively. The fluctuations of intensity within regions of the same sign suggest these modes may represent a planar slice of out-of-plane structures. Despite this, these modes have energy shares of $23.67\%$ and $21.96\%$, respectively, and thus represent almost half of the reduced-order dynamics of this sphere.
@@ 4 Discussion
The first four DMD modes of each sphere, including the SHS sphere, represent different instability scales in the shear layer of the wake. This differs from the leading POD modes in [40], where the first six SHS modes represented multiple permutations of the structure equivalent to the first two POD modes of the untreated spheres. This was likely due to the interaction between the wake and the plastron resulting in a greater degree of three-dimensional variation in the wake, as observed in settling sphere experiments [14]. Thus, the comparison of these modes allows for an examination of the effect of the superhydrophobic surface treatment on the shear layer instabilities. The wavelengths of these modes are consistently longer for the PRS case than the REF case, and their decay rates are equivalent or slower. Additionally, while the modes of the REF sphere were consistently symmetrical, the symmetry was reduced in the PRS case and further diminished in the SHS case for some modes. The structures near the sphere in the first PRS mode are closer to the centreline than those in the first REF mode, returning to a similar radial position further downstream. This may be due to a tripping effect from the pores, similar to the effect of dimples on a golf ball [52], delaying separation. The effect of the superhydrophobic surface treatment on the shear layer is more varied. The second SHS mode has a significantly shorter wavelength and a faster decay rate than the corresponding modes of the untreated spheres. The mode shape is also significantly altered compared to the untreated spheres, in which the structures were mostly horizontally aligned. The SHS mode instead features structures which overlap in the streamwise direction, with the structures extending downstream and away from the centreline to overlap with the portion of the upstream structures close to the centreline. The first SHS mode, which is the closest equivalent to the second modes of the untreated spheres, has a longer wavelength and faster decay rate. The SHS mode also features additional small structures above the regions which correspond to those present in the untreated sphere modes. The larger structures in this mode are also closer to the centreline than those in the modes of the untreated spheres, which may result from the slip introduced by the surface treatment delaying flow separation. The changes to the flow over the SHS sphere are likely limited to the near-wake region, as the structures in these modes approach shapes similar to those of the modes of the untreated spheres towards the end of the measurement domain. Without measurements at the surface of the sphere, which were not possible for this study due the laser reflections through the plastron, the exact cause of these changes cannot be determined. However, the likely candidates are the introduction of slip delaying separation until further back on the sphere, the change in viscosity from the plastron dissipating or the injection of micro-bubbles into the wake.
The third and fourth modes of the PRS sphere have longer wavelengths and decay more slowly than those of the REF modes, with the SHS modes having even longer wavelengths and slower decay rates than those of the PRS sphere. However, the change in the structure of these modes is less significant than those observed in the first and second modes. The shapes of these modes are comparable across the three spheres, with the differences in wavelength being the only discernible variation.
The fifth modes of the untreated spheres vary significantly in shape from the first four modes, although they remain similar to each other. As the structures in this mode extend through the centreline and change form around $x/D=3$, these modes may represent interactions between the recirculation region and the shear layers on either side of the sphere. However, because the present study is focused on the shear layer and these modes clearly extend beyond the domain, the insight provided by these modes is limited.
The sixth modes of the untreated spheres, as well as the fifth and sixth modes of the SHS sphere, had wavelengths longer than could be resolved within the measurement domain. Despite this, the mode shapes for the untreated spheres loosely resembled the structures in the leading four modes, but extended significantly in the streamwise direction. These modes also had the largest energy share for their respective spheres, consistent with the energy of the other modes decreasing with higher wavenumbers. In the SHS case, both the fifth and sixth modes, which made up almost half of the energy share of the reduced-order dynamics, had wavelengths significantly longer than the domain. This is similar to the shift mode represented by the DMD of a transient cylinder wake presented in [53], and may reflect the spatial decay of turbulence in the streamwise direction. The presence of two of these modes for the SHS case suggests that this decay is more complex than the untreated sphere cases, perhaps due to the interaction between the plastron and the flow. The smaller structures visible in these modes may also suggest an effect on the turbulent fluctuations from the plastron dissolving into the flow at a regular rate.
Despite these limitations, the current analysis was able to elucidate modes representing the instabilities in the shear layer, enabling an assessment of the effect of the superhydrophobic surface treatment. In contrast to the POD analysis on the same spheres, the DMD applied in the current work was able to identify changes in the flow due to the addition of pores. The DMD modes of the SHS were not affected by the three-dimensionality of the dominant instability in the wake, which dominated the leading POD modes in [40]. Thus, the leading modes of the SHS sphere were able to clearly capture the equivalent instabilities to the untreated spheres. Additionally, the fifth and sixth modes of the SHS sphere captured smaller structures within their main structure, which were not present in the equivalent modes of the untreated spheres.